A base diameter is 14cm while the height is 12cm. Calculate the total surface area if the cylinder has both a base and a top (π=22/7)


1
Expert's answer
2021-06-14T15:46:42-0400

Let "d=" a base diameter, "h=" the height of the cylinder. Then


"Area=A=2\pi(\dfrac{d}{2})^2+2\pi(\dfrac{d}{2})h=\dfrac{\pi d}{2}(d+2h)"


"A=\dfrac{\dfrac{22}{7}(14) }{2}(14+2(12))=836 (cm^2)"

The total surface area if the cylinder is "836" cm2.